OpenStudy (anonymous):

What is the maximum angle at which you can throw a ball so that its distance from you never decreases during its flight?

OpenStudy (mathmate):

There may be more elaborate methods, but the following should work: Use kinematic equations: u=initial velocity (m/s) \(a=initial angle with horizontal (deg)\) t=elapsed time (s) g=-9.8 m/s^2 = acceleration due to gravity then x=u cos(a)t y=u sin(a)t+(1/2)gt^2 Distance from original position d(t)=sqrt(x^2+y^2) Square of distance with respect to time D(t) = d^2=x^2+y^2 Evaluate D(t) and find "a" such that D'(t) is increasing for t>0.